# Image Inpainting based on Visual-Neural-Inspired Specific Object-of-Interest Imaging Technology

### v2

@@ Abstract
Conventional image inpainting methods operate on whole images. Drawing on cortical processing principles this study addresses the bottlenecks of conventional holistic image inpainting methods–susceptibility to informational redundancy and low computational efficiency under occlusions and complex backgrounds, by proposing a novel framework: ”Specific Object-of-Interest Imaging”. This stage extracts and encodes object-level representations from complex scenes, generating semantic and structural priors that can be seamlessly integrated into any inpainting framework. On our Teapot object dataset, Elephant object dataset, Giraffe object dataset and Zebra object dataset experimental validation demonstrates–compared with the repair model without using this method, the repair model using this method has more advantages, across metrics including SSIM, PSNR, MAE, and LPIPS, while maintaining robustness in extreme scenarios (low illumination, high noise, multi-object occlusion, motion blur). Theoretical analysis integrated with cognitive neuroscience perspectives reveals profound correlations between the ”object precedence perception” mechanism and dynamic feature modulation in visual cortices (V1–V4). We demonstrate that incorporating Stage I significantly enhances object consistency and semantic coherence across various mainstream inpainting models. This work not only achieves efficient and precise target-centric imaging but also pioneers interdisciplinary pathways bridging brain-inspired computational frameworks with advanced image inpainting techniques.
@@
organization=School of Electronic Science and Engineering (School of Microelectronics), South China Normal University, addressline=Foshan, 528000, country=China
@@
organization=St. Petersburg School of Shipbuilding and Ocean Technology, Guangdong Ocean University, addressline=Zhanjiang 524088, country=China
@@
organization=Robotics Laboratory, Institute of Automation and Control Processes, addressline=Far Eastern Branch of the Russian Academy of Sciences, city=Vladivostok, postcode=690041, country=Russia
[…]
organization=Robotic Laboratory, Institute of Automation and Control Processes FEB RAS, city=Vladivostok, postcode=690041, country=Russia
@@ 1 Introduction
Image Inpainting, a foundational and critically important task in computer vision and image processing, aims to restore missing or corrupted regions. Such impairments may arise from accidental occlusions during capture, sensor-induced noise, data loss in encoding/decoding or transmission, intentional object removal in post-processing, or physical deterioration in historical photographs. The capacity to restore image integrity, visual authenticity, and semantic plausibility renders this technology indispensable across diverse applications, including cultural heritage preservation, digital photography enhancement, medical image reconstruction, cinematic visual effects production, and the interpretation of occluded scenarios in autonomous driving systems. However, producing highly accurate and semantically coherent inpainting results remains a fundamental challenge.
@@
Conventional inpainting methods primarily rely on intrinsic statistical properties and local structural information of images [1, 2, 3, 4, 5, 6, 7, 8]. Canonical approaches include diffusion- and patch-based algorithms, whose core principle involves leveraging neighborhood information from intact regions to fill missing areas via propagation or copy-paste mechanisms. While adequate for small-scale, texturally regular regions, these methods exhibit significant limitations when confronted with extensive missing areas, complex structural boundaries, or semantically rich regions. They often generate artifacts such as blurred textures, fractured structural continuity, and semantically inconsistent content that undermines global image context. These limitations significantly constrain their effectiveness in real-world, high-fidelity imaging applications.
@@
The rapid development of deep learning—particularly Convolutional Neural Networks (CNNs), Generative Adversarial Networks (GANs), and emerging Transformer- and Diffusion-based models—has driven breakthroughs in complex pattern learning and high-fidelity image synthesis, thereby catalyzing a paradigm shift in image inpainting [9, 10, 11, 12, 13, 14, 15]. These data-driven approaches capture rich image priors and deep semantic representations from large-scale datasets, thereby enabling models to predict missing content with improved accuracy. These models reconstruct intricate textural details while maintaining plausibility. For example, restored objects exhibit geometrically consistent structures, physically coherent illumination, and contextually appropriate depth relationships. State-of-the-art models can generate diverse yet equally plausible results, demonstrating performance that substantially surpasses conventional methods.
@@
Despite these advances and their widespread deployment, researchers still face persistent and critical challenges. A pivotal yet often overlooked issue is the inefficiency and inaccuracy stemming from informational redundancy. In many real-world applications, inpainting is fundamentally object-centric rather than indiscriminately processing the entire frame. Under such conditions, prevailing whole-image inpainting models — encompassing nearly all contemporary deep learning methods — incur substantial computational overhead by processing redundant background information. For instance, autonomous vehicles equipped with 8K cameras can generate up to 1.5 TB/s of data, with nearly 90% consisting of redundant background or irrelevant content; military reconnaissance satellites transmit massive raw imagery, with less than 0.1% containing actionable intelligence; and in medical X-ray imaging and industrial non-destructive testing, conventional area-array detectors achieve less than 0.13% efficiency in extracting useful information. Whole-image models consume substantial computational resources analyzing redundant background information, leading to severe resource wastage and reduced accuracy in target-object reconstruction due to background interference, calling for a paradigm shift from holistic processing to targeted, object-focused inference.
In confronting these challenges, biological vision systems offer an elegant and efficient solution. A core component of biological vision is the attention mechanism, which prioritizes task-relevant targets while suppressing redundant background information. This enables efficient, accurate, and energy-optimized information processing for target recognition and imaging in cluttered or dynamic scenes. This “object-first perception” strategy offers a promising solution for addressing low information density and background noise in inpainting: an inpainting pipeline could first localize and model target-object structures while excluding irrelevant background as interference. Such an approach would enhance imaging focus, computational efficiency, and output quality.
Building on this analysis of the current state of inpainting, its technical bottlenecks, and biological inspiration, we propose a novel front-end paradigm: Image Inpainting via Visual-Neural-Inspired Specific Object-of-Interest Imaging technology(SIOI). Inspired by biological “object-first perception”, we propose a paradigm shift: Specific Object-of-Interest Imaging (SIOI). Unlike traditional methods, SIOI decouples the ‘what’ and ‘where’ of the object from the ‘how’ of its inpainting. It acts as a generic front-end that extracts a clean object representation, suppressing background noise, and can be plugged into any existing inpainting model. Our framework decomposes the imaging process into two distinct stages. Stage 1 focuses exclusively on extracting and modeling the structural information of target objects. This stage constitutes a dedicated information-extraction process designed to maximize target-relevant features while minimizing or eliminating background interference. Stage 2 performs coarse structural restoration guided by refined priors to support detail synthesis, refines the restoration, ensuring visual coherence and semantic consistency in the final output. The key innovation lies in Stage I; Stage II is deliberately kept simple in our experiments to prove the concept. By frontloading SIOI, the pipeline significantly reduces background distractions and concentrates structural modeling on critical target regions. This substantially reduces the data volume for processing, thereby improving computational efficiency while enhancing structural continuity and semantic accuracy. As shown in Tab. 1, unlike other image-restoration methods, this bio-inspired approach aligns with the hierarchical modeling of the visual cortex in neural networks, ensuring that computational efficiency parallels human attentional selection.
@@
    We introduce Specific Object-of-Interest Imaging (SIOI), a novel front-end paradigm for object-centric image processing. SIOI explicitly decouples target-object extraction from texture synthesis, providing a generic mechanism to reduce background interference and generate powerful structural priors for any subsequent task.
@@
    We conduct a preliminary exploration of correlations between human vision and deep networks, integrating neuroscientific and psychological insights on hierarchical visual cortex processing to establish theoretical foundations.
@@
    We demonstrate that SIOI is an effective, plug-and-play front-end module. Through extensive experiments, we show that integrating SIOI consistently enhances the performance of various mainstream inpainting models across multiple metrics and datasets, while significantly improving robustness in extreme scenarios.
@@
Traditional Digital image processing and classical machine learning techniques Handcrafted features and algorithms Entire image Not considered
Deep-Learning Data-driven approaches enabled by deep neural networks Learned feature representations Entire image Not considered
Ours Bio-inspired two-stage framework Object perception → Object inpainting Specific Object-of-Interest V1-V4 hierarchical simulation
@@ 2 Related Works
Image inpainting is a long-studied and widely applied problem whose methods have undergone substantial paradigm shifts. From early mathematical models to modern deep generative architectures. While these approaches differ in implementation, they overwhelmingly share a whole-image processing paradigm.
@@ 2.1 Traditional Mathematical Algorithms: Digital Image Processing and Machine Learning
This period emphasized leveraging low-level statistical properties, local geometric structures, and hand-crafted mathematical models. Representative approaches include: 1) PDE- and diffusion-based models [1, 16, 17], which use isophote directions or engineered diffusion tensors to propagate color and gradient information from intact regions into missing areas while preserving edge coherence; and 2) patch-based methods [18, 19, 20], which exploit non-local self-similarity by searching for patches in intact regions that best match neighborhoods around missing regions and iteratively copying them to fill holes. Traditional mathematical approaches [21, 22, 23] reconstruct missing areas by propagating local structures or copying similar patches. They perform reasonably on small, regular textures but break down under large occlusions or semantically rich regions, since they lack object-level understanding.
@@ 2.2 Data-Driven Neural Network Approaches
Deep-learning approaches substantially improve structural coherence and realism by learning high-level priors. Predominant paradigms include: (1) CNN encoder–decoder architectures [24, 4, 25, 26, 27], which exploit multi-level encoder features to progressively generate missing content in decoders, attention mechanisms and gated convolutions further improve focus on relevant regions; (2) GAN-based methods [28, 29, 30, 31, 32, 33, 34], which employ adversarial training between generators and discriminators to produce highly realistic outputs and excel at fine texture synthesis and alignment with natural image distributions; (3) Transformer-based approaches [4, 35, 36, 37, 38, 39, 40], which leverage self- and cross-attention to model long-range dependencies between patches and thereby capture global semantics and structural consistency, improving performance on complex scenes and large corruptions; and (4) diffusion models [41, 42, 43], which cast generation as a progressive denoising process that iteratively recovers image details from noise, offering fine-grained detail control, diverse outputs, and high fidelity.
Yet, most still operate on the entire image, wasting computation on irrelevant background and suffering from interference when target and background overlap. Crucially, they rarely encode explicit object guidance, limiting their ability to ensure object consistency in inpainting. Whole-image models thus suffer from several issues in object-centric tasks: 1) information redundancy and computational inefficiency — processing the entire image wastes computation on irrelevant background content, exacerbating the low effective information density typical in domains such as autonomous driving and satellite imaging; 2) background interference — complex or misleading backgrounds can corrupt the reconstruction of target object structure and semantics, especially when target and background share similar textures or colors; and 3) lack of target-specific guidance and focus — current methods rarely model explicitly which object should be targeted and therefore struggle to separate target signals from background noise during inpainting. Although some methods adopt user sketches or bounding boxes as guidance, these require extra interaction and still struggle to achieve precise target separation in complex scenes.
Thus, a biologically inspired, target-focused imaging paradigm — one that actively suppresses irrelevant background information — is imperative for object-centric tasks. Our proposed method, Inpainting via Specific Object-of-Interest Imaging , directly addresses this gap. It introduces a front-loaded Specific Object-of-Interest Imaging module that algorithmically maximizes target-relevant information while suppressing background interference. This yields highly refined, target-centric signals for subsequent inpainting stages and thereby mitigates information redundancy and background interference.
@@ 2.3 Insights from Human Visual Mechanisms for Image imaging
Psychologists and neuroscientists have reported direct correlations between deep neural networks (DNNs) and hierarchical processing in the visual cortex [44, 45]. Human vision follows a hierarchical pathway: retina $\rightarrow$ lateral geniculate nucleus (LGN) $\rightarrow$ V1 (edge extraction) $\rightarrow$ V4 $\rightarrow$ IT (semantic processing) [46, 47]. This observation motivates multi-stage imaging networks [48, 49], in which early convolutional layers focus on local structure repair (analogous to V1 edge detection), while deeper layers integrate object identity and scene context (analogous to IT semantic processing); skip connections can be viewed as a crude analogue of cortical feedback. Predictive-coding theory [50, 51] proposes that the brain leverages prior knowledge (e.g., Gestalt principles) to generate predictions and correct sensory errors. This idea parallels the generator–discriminator framework in deep inpainting [24, 4]: generators predict missing content by learning data distributions, while discriminators enforce semantic plausibility via adversarial training [52]. At the level of attentional regulation, bottom-up attention captures salient features (e.g., high-frequency details) via thalamic structures such as the pulvinar [53, 54]; this motivates algorithmic mechanisms such as saliency-map-weighted losses that emphasize abrupt edge regions [55]. Conversely, top-down attention involves prefrontal-cortex (PFC) modulation of visual responses according to task goals [54]. This inspires task-oriented strategies in models: for example, conditional vectors can steer generation to enhance discriminative feature repair (e.g., facial keypoint reconstruction) and to improve global structural consistency for scene reconstruction. Examples of biologically inspired techniques include contour interpolation via edge-consistency losses [56], motivated by V1/V2 responses to illusory contours [57, 58, 59]. Semantic guidance has been realized by mapping IT-like object activations. For instance, CLIP-guided diffusion has been used to align textual concepts with image generation [42]. Memory-augmented imaging methods simulate hippocampal–cortical episodic recall by retrieving reference segments from external memory banks [60]. Consequently, advancing image inpainting benefits from drawing principled inspiration from human visual mechanisms to identify promising directions for improvement. This ‘object-first’ perception allows efficient suppression of background noise and allocation of resources to relevant targets. Such mechanisms directly motivate our proposed Specific Object-of-Interest Imaging, which introduces an analogous front-end object-first module into the inpainting pipeline.”
In summary, while existing methods have advanced in realism and fidelity, they all share a whole-image bias that is ill-suited for object-centric tasks. Inspired by biological vision, we propose a new front-end paradigm—Specific Object-of-Interest Imaging—to fill this critical gap.
@@ 3 Proposed Methodology
This section delineates our proposed framework for Image Inpainting. Our approach introduces a novel paradigm for this task. On a conceptual level, it shares a similar goal—focusing on a specific object region—with other imaging techniques, such as [61, 62, 63, 64, 65], but the underlying methodology is uniquely our own work. The processing pipeline is illustrated in Fig. 1.
@@ 3.1 Specific Object-of-Interest Imaging
Specific Object-of-Interest Imaging is also known as Attention-Targeted Imaging. Desimone and Duncan’s competitive-bias model [66] posits that multiple objects compete for activation resources in the visual cortex. Attention serves as a bias signal that preferentially amplifies target objects and suppresses background information. This provides a neuroscientific foundation for implementing background suppression and object-enhancement saliency selection in computational networks. O’Craven et al. [67] demonstrated with fMRI that objects act as fundamental attentional units, exhibiting holistic enhancement even when component features are spatially segregated. Other studies [68, 69] further showed that object boundaries automatically attract attention, recruiting V1–V4 regions and intensifying internal feature encoding. Perry et al. [70] reported that foveal cortical magnification increases encoding density, and that task-dependent plasticity enables high-resolution perception of behaviorally relevant regions. Carrasco and Barbot [71, 72] systematically characterized attention’s dual regulation of spatial resolution: exogenous attention enhances frequency sensitivity at fixation loci, whereas endogenous attention reallocates resolution according to task demands. Barbot and Carrasco further showed that attentional guidance amplifies high-frequency recognition, providing a basis for region-adaptive reconstruction. Ekman et al. [73] demonstrated that attentional signals propagate along object boundaries in V1, inducing internal feature enhancement — a phenomenon translatable to frequency-domain focusing mechanisms in networks (e.g., concentrating computational resources using object masks).Collectively, these findings reveal an object-based prioritization mechanism: the visual system preferentially allocates resources to enhance detailed representation of target objects. These observations directly inspired our object-detection–guided saliency-enhancement branch, which applies high-frequency component weighting to target regions. We propose an approach that integrates saliency modeling with biologically inspired, object-guided image imaging to construct a cognitively plausible, interpretable, and task-relevant image inpainting system (see Fig. 2).
@@
Neuroanatomically, visual processing follows a hierarchical architecture: local feature detection (retina) $\rightarrow$ spatial integration (LGN) $\rightarrow$ orientation selectivity (V1) $\rightarrow$ holistic object representation (IT cortex). This multi-stage processing implies that a natural-scene image $k$ can be viewed as a superposition of multiple latent target objects, with attentional modulation in V4 amplifying target-specific features while suppressing background information. Grounded in neuronal population-coding theory and Hebbian plasticity, we formalize this composition as a linear superposition model:
@@
where $k\in\mathbb{R}^{H\times W\times C}$ denote the scene tensor, $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ the $i$-th target object, and $M\in\mathbb{Z}^{+}$ the total number of objects. Each $k_{i}$ corresponds to a semantically complete entity in the scene (e.g., pedestrians or vehicles). Amodal content completion emphasizes restoring full object structures in occluded regions. The core objective is anatomically precise object separation: accurately extracting specific targets $k_{i}$ ($i\in\{1,\dots,M\}$) from complex backgrounds. This parallels the “cocktail-party” effect in human perception when attending to a target in cluttered scenes.
To overcome the representational fragility of traditional methods in noisy environments, we propose a biologically inspired dynamic feature-modulation paradigm. Its mathematical core is a target-sensitive measurement matrix $\Gamma$. Via a differentiable learning mechanism, $\Gamma$ induces a cross-scale feature set $\{\xi_{j}\}_{j=1}^{N}$, reflecting the hierarchical abstraction from V1 simple cells (edge/texture responses) to IT complex cells (holistic shape responses). Below we make these notions precise.
For a target index $\mu\in\{1,\dots,M\}$, we write
@@
where $\Gamma\in\mathbb{R}^{N\times D}$ denotes the measurement matrix, $\xi_{j}\in\mathbb{R}^{D}$ the feature-basis vectors, and $\omega_{j,\mu}\in\mathbb{R}$ the feature weights associated with target $\mu$. When optimization drives the object- and target-specific weights $\omega_{j,\mu}$ to concentrate on the feature patterns of a particular $\mu$, the system exhibits a selective-gain effect: scene measurements become dominated by target-related information flow. These dynamics are expressed as
$\Gamma k\;=\;\begin{bmatrix}\sum_{m=1}^{M}\eta_{1,m}\,\xi_{1}\\[4.0pt] \vdots\\[4.0pt] \sum_{m=1}^{M}\eta_{N,m}\,\xi_{N}\end{bmatrix}\approx\widehat{k}_{\mu},$ (3)
where $\eta_{j,m}$ denotes the contribution of object $m$ to feature channel $j$. Measurements concentrate on the specific target when $\eta_{j,m}\to\delta_{m\mu}$ for each feature channel $j$, i.e., when the contributions from non-target objects vanish.
Functionally, this formulation mirrors the hierarchical transformation from V1 simple cells (edge/texture responses) to IT complex cells (holistic shape responses), with $\Gamma$ playing the role of attentional gain control in V4. V1 encodes local orientations, while V4 integrates global contours under top-down modulation, consistent with our FFC-based frequency-domain focusing. From a functional-analysis perspective, this process can be interpreted as a target-guided modulation on an appropriate Riemannian manifold:
@@
where $\otimes$ denotes the diffeomorphism operator in the feature space. Establishing a rigorous mathematical model is essential for precise object separation. Consider an input scene containing $N$ latent targets. Sparse representation of individual objects. Each object $k_{i}$ ($i=0,1,\dots,N-1$) can be represented sparsely in a high-dimensional dictionary basis $\Psi$:
@@
where $\psi_{ij}$ denotes the adaptive sparse basis and $x_{ij}$ the sparse coefficients. Panoramic scene decomposition indicates that the holistic scene can be expressed as a linear combination of objects:
@@
subject to the following constraints: $\|\mathbf{x}_{i}\|_{0}\ll M\quad\text{(Sparsity)},\quad\psi_{ij}\in\mathbb{R}^{D}\quad\text{(Dictionary Basis)},\quad x_{ij}\in\mathbb{R}\quad\text{(Sparse Coefficient)}.$ This representation satisfies the principles of compressed sensing, thereby ensuring object separability in the feature space. This theoretical model is implemented through a bio-inspired neural architecture. The Feedforward Feature Selection Network simulates the ventral (“what”) visual pathway by performing target-sensitive filtering on masked defective inputs $M_{h}$. The Target Reconstruction Network mimics the spatial integration function of the parietal cortex and is dedicated to target reconstruction. The training process is formulated as a physically constrained optimization problem:
@@
where $\mathcal{C}(\cdot)$ denotes the loss function. Upon convergence, the FECNet weights $\xi$ encode target knowledge into a feature filter $\phi$. Applying $\phi$ to the panoramic scene yields:
@@
Since $\phi$ is a linear operator, it satisfies the following projection property:

### v3

# Visual-Neural-Inspired Image Inpainting for Specific Objects-of-Interest Imaging

@@ Abstract
Conventional image inpainting techniques typically process entire images, which often leads to computational inefficiency and susceptibility to information redundancy, particularly in occluded or cluttered scenes. Inspired by cortical processing mechanisms, this study introduces a novel framework termed “Specific Object-of-Interest Imaging” (SIOI) to overcome these limitations. The proposed approach first extracts and encodes object-level representations from complex scenes, producing structural and semantic priors that can be seamlessly integrated into existing inpainting pipelines. Extensive evaluations on dedicated object datasets—Teapot, Elephant, Giraffe, and Zebra—demonstrate that models incorporating SIOI consistently outperform those without it across key metrics including SSIM, PSNR, MAE, and LPIPS. The framework also exhibits strong robustness under challenging conditions such as low illumination, high noise, multiple object occlusions, and motion blur. Furthermore, theoretical analysis grounded in cognitive neuroscience reveals meaningful connections between the “object-first” perceptual mechanism and dynamic feature modulation in visual cortical areas (V1–V4). By significantly enhancing object consistency and semantic coherence across diverse inpainting models, this work not only advances efficient and target-aware image restoration but also fosters interdisciplinary convergence between brain-inspired computation and cutting-edge inpainting technologies.
@@
[1]organization=School of Electronic Science and Engineering (School of Microelectronics), South China Normal University, addressline=Foshan, 528000, country=China
@@
[2]organization=St. Petersburg School of Shipbuilding and Ocean Technology, Guangdong Ocean University, addressline=Zhanjiang 524088, country=China
@@
[3]organization=Robotics Laboratory, Institute of Automation and Control Processes, addressline=Far Eastern Branch of the Russian Academy of Sciences, city=Vladivostok, postcode=690041, country=Russia
@@ 1 Introduction
Image inpainting is a fundamental computer vision task that restores missing or corrupted image regions. These imperfections can result from occlusions, sensor noise, data transmission loss, intentional editing, or physical damage. This capability makes it vital for applications like cultural preservation, photography, medical imaging, visual effects, and autonomous systems. However, achieving accurate and semantically coherent inpainting remains challenging.
@@
Conventional inpainting methods primarily rely on intrinsic statistical properties and local structural information of images [1, 2, 3, 4, 5, 6, 7, 8], whose core principle involves leveraging neighborhood information from intact regions to fill missing areas via propagation or copy-paste mechanisms. While adequate for small-scale, texturally regular regions, these methods exhibit significant limitations when confronted with extensive missing areas, complex structural boundaries, or semantically rich regions. They often generate artifacts such as blurred textures, fractured structural continuity, and semantically inconsistent content that undermines global image context.
@@
The rapid development of deep learning—particularly Convolutional Neural Networks (CNNs), Generative Adversarial Networks (GANs), and emerging Transformer- and Diffusion-based models—has driven breakthroughs in complex pattern learning and high-fidelity image synthesis, thereby catalyzing a paradigm shift in image inpainting [9, 10, 11, 12, 13, 14, 15]. These data-driven approaches capture rich image priors and deep semantic representations from large-scale datasets, thereby enabling models to predict missing content with improved accuracy. These models reconstruct intricate textural details while maintaining plausibility.
@@
Despite these advances, significant challenges remain. A key overlooked issue is inefficiency from information redundancy. In practice, inpainting is often object-centric rather than holistic. Consequently, most existing methods process redundant background, incurring substantial overhead. For example, across applications like autonomous driving, reconnaissance, and medical imaging, typically over 90% of data is background rather than target information. Whole-image models consume substantial computational resources analyzing redundant background information, leading to severe resource wastage and reduced accuracy in target-object reconstruction due to background interference, calling for a paradigm shift from holistic processing to targeted, object-focused inference.
In confronting these challenges, biological vision offers an efficient solution. Visual attention mechanisms prioritize relevant targets while suppressing background, enabling efficient target recognition in complex scenes. This ”object-first” approach addresses information sparsity by first localizing targets while excluding background interference. This improves focus, efficiency, and output quality.
Based on this analysis, we propose Specific Object-of-Interest Imaging (SIOI). SIOI decouples object identification from inpainting execution. It serves as a generic front-end that extracts clean object representations while suppressing background, pluggable into any inpainting model. Our framework employs two stages: (1) object perception, and (2) object inpainting. SIOI reduces background interference and focuses processing on target regions, improving efficiency and accuracy. As shown in Tab. 1, this approach aligns with visual cortex hierarchy, mimicking human attentional mechanisms.
@@
    We propose Specific Object-of-Interest Imaging (SIOI), a novel framework that decouples object extraction from texture synthesis to reduce background interference and generate structural priors for subsequent tasks.
@@
    We investigate correlations between human vision and deep networks, integrating neuroscientific insights from visual cortex hierarchy to establish theoretical foundations.
@@
    We demonstrate SIOI’s effectiveness as a plug-and-play module through comprehensive experiments, showing consistent performance improvements across metrics, datasets, and extreme scenarios.
@@
Traditional Classical image processing Handcrafted features Entire image Not considered
Deep-Learning Deep neural networks Learned representations Entire image Not considered
Ours Bio-inspired Object perception → inpainting Specific Object-of-Interest V1-V4 simulation
@@ 2 Related Works
Image inpainting has evolved through substantial paradigm shifts, from early mathematical models to modern deep generative architectures. While these approaches differ in implementation, these methods predominantly employ whole-image processing.
@@ 2.1 Traditional Mathematical Algorithms: Digital Image Processing and Machine Learning
This period emphasized low-level statistics, local geometric structures, and hand-crafted models. Representative approaches include: 1) PDE- and diffusion-based models [1, 16, 17] that propagate color and gradient information while preserving edges; and 2) patch-based methods [18, 19, 20] that exploit non-local self-similarity by matching and copying patches from intact regions. These approaches [21, 22, 23] reconstruct missing areas through local propagation or patch copying. They perform well on small regular textures but struggle with large occlusions or semantically rich regions due to limited object-level understanding.
@@ 2.2 Data-Driven Neural Network Approaches
Deep learning methods improve structural coherence and realism by learning high-level priors. Key paradigms include: (1) CNN encoder-decoder architectures [24, 4, 25, 26, 27] using multi-level features to generate content, with attention mechanisms enhancing focus; (2) GAN-based approaches [28, 29, 30, 31, 32, 33, 34] using adversarial training to produce realistic outputs and fine textures; (3) Transformer-based methods [4, 35, 36, 37, 38, 39, 40] leveraging self- and cross-attention to model long-range dependencies and capture global semantics; and (4) Diffusion models [41, 42, 43] employing progressive denoising to recover image details with fine-grained control and high fidelity.
Yet, most still operate on the entire image, wasting computation on irrelevant background and suffering from interference when target and background overlap. Crucially, they rarely encode explicit object guidance, limiting their ability to ensure object consistency in inpainting. Whole-image models thus suffer from several issues in object-centric tasks: 1) information redundancy and computational inefficiency — processing the entire image wastes computation on irrelevant background content, exacerbating the low effective information density typical in domains such as autonomous driving and satellite imaging; 2) background interference — complex or misleading backgrounds can corrupt the reconstruction of target object structure and semantics, especially when target and background share similar textures or colors; and 3) lack of target-specific guidance and focus — current methods rarely model explicitly which object should be targeted and therefore struggle to separate target signals from background noise during inpainting. Although some methods adopt user sketches or bounding boxes as guidance[44, 45, 46], these require extra interaction and still struggle to achieve precise target separation in complex scenes.
Therefore, a biologically inspired, target-focused imaging paradigm that actively suppresses irrelevant background information is essential for object-centric tasks. Our method, Inpainting via Specific Object-of-Interest Imaging, directly addresses this gap. It introduces a front-end Specific Object-of-Interest Imaging module that algorithmically maximizes target-relevant information and suppresses background interference. This produces refined target-centric signals for subsequent stages, mitigating information redundancy and background interference.
@@ 2.3 Insights from Human Visual Mechanisms for Image imaging
[…]
In summary, despite advances in realism, existing methods retain a whole-image bias ill-suited for object-centric tasks. Inspired by biological vision, we propose Specific Object-of-Interest Imaging to address this gap.
@@ 3 Proposed Methodology
This section presents our novel image inpainting framework. While conceptually sharing the object-focus goal with other techniques [63, 64, 65, 66, 67], our methodology is fundamentally distinct. The processing pipeline is illustrated in Fig. 1.
@@ 3.1 Specific Object-of-Interest Imaging
Also termed Attention-Targeted Imaging, SIOI builds on neuroscientific foundations. Desimone and Duncan’s model [68] posits visual cortex competition among objects for activation resources. Attention acts as a bias signal amplifying targets while suppressing background, providing a neuroscientific basis for computational background suppression and object enhancement. O’Craven et al. [69] fMRI studies show objects as fundamental attentional units with holistic enhancement. Studies [70, 71] show object boundaries automatically attract attention, engaging V1-V4 regions. Perry et al. [72] reported foveal magnification increases encoding density and task-dependent plasticity enables high-resolution perception. Carrasco and Barbot [73, 74] characterized attention’s dual regulation: exogenous enhances frequency sensitivity, endogenous reallocates resolution. They further showed attentional guidance amplifies high-frequency recognition, enabling region-adaptive reconstruction. Ekman et al. [75] showed attentional signals propagate along V1 object boundaries, inducing feature enhancement—translating to network frequency-domain focusing. Collectively, these reveal object-based prioritization: preferential resource allocation to enhance target representation. These directly inspired our object-detection-guided saliency enhancement applying high-frequency weighting to targets. We integrate saliency modeling with bio-inspired object guidance to construct a cognitively plausible inpainting system (Fig. 2).
@@
Anatomically, the visual system processes information hierarchically: local feature detection (retina) $\rightarrow$ spatial integration (LGN) $\rightarrow$ orientation selectivity (V1) $\rightarrow$ holistic object representation (IT cortex). This hierarchical processing suggests that a natural image $k$ comprises multiple latent objects, where V4-mediated attentional modulation enhances target features and suppresses background information. Based on neuronal population-coding and Hebbian plasticity, we model this composition through linear superposition:
@@
where $k\in\mathbb{R}^{H\times W\times C}$ is the scene tensor, $k_{i}\in\mathbb{R}^{h_{i}\times w_{i}\times C}$ represents the $i$-th target object, and $M\in\mathbb{Z}^{+}$ is the total object count. Each $k_{i}$ represents a semantically coherent entity (e.g., a pedestrian or vehicle). Amodal completion focuses on reconstructing fully occluded object structures. The core objective is precise object separation—extracting specific targets $k_{i}$ ($i\in\{1,\dots,M\}$) from cluttered backgrounds. This mirrors the “cocktail-party” effect, where humans focus on a target speaker amidst noise.
To overcome the limitations of traditional methods in noise, we introduce a bio-inspired dynamic feature modulation approach. Its mathematical basis is a target-sensitive measurement matrix $\Gamma$. Through differentiable learning, $\Gamma$ produces a multi-scale feature set $\{\xi_{j}\}_{j=1}^{N}$, capturing the hierarchy from V1 simple cells (edge/texture) to IT complex cells (shape representation). We formalize these concepts next.
For a target index $\mu\in\{1,\dots,M\}$, we have:
@@
where $\Gamma\in\mathbb{R}^{N\times D}$ is the measurement matrix, $\xi_{j}\in\mathbb{R}^{D}$ are the basis vectors, and $\omega_{j,\mu}\in\mathbb{R}$ are the target-specific weights. When optimized, the weights $\omega_{j,\mu}$ concentrate on patterns for $\mu$, creating a selective-gain effect where scene measurements are dominated by the target. This leads to the scene-level measurement:
$\Gamma k=\begin{bmatrix}\sum_{m=1}^{M}\eta_{1,m}\xi_{1}\\ \vdots\\ \sum_{m=1}^{M}\eta_{N,m}\xi_{N}\end{bmatrix}\approx\widehat{k}_{\mu},$ (3)
where $\eta_{j,m}$ is the contribution of object $m$ to channel $j$. Measurements focus on the target when $\eta_{j,m}\to\delta_{m\mu}$ for all $j$, meaning non-target contributions vanish.
Functionally, this models the transformation from V1 (edges/textures) to IT (shapes), with $\Gamma$ acting as an attentional gain in V4. V1 encodes local orientations, while V4 integrates global contours, aligning with our frequency-domain (FFC) focus. From a functional view, this process is a target-guided modulation on a Riemannian manifold:
@@
where $\otimes$ is the diffeomorphism operator in feature space. A rigorous mathematical model is essential for precise object separation. For a scene with $N$ latent targets, each object admits a sparse representation. Each object $k_{i}$ ($i=0,1,\dots,N-1$) is sparsely represented in a high-dimensional basis $\Psi$:
@@
where $\psi_{ij}$ is the adaptive basis and $x_{ij}$ the sparse coefficients. The scene decomposes into a linear combination of objects:
@@
subject to sparsity ($\|\mathbf{x}_{i}\|_{0}\ll M$), basis constraints ($\psi_{ij}\in\mathbb{R}^{D}$), and coefficient constraints ($x_{ij}\in\mathbb{R}$). This compressed sensing formulation ensures object separability in feature space.
We implement this via a bio-inspired neural architecture. The Feedforward Feature Selection Network mimics the ventral (”what”) pathway, filtering masked inputs $M_{h}$. The Target Reconstruction Network emulates parietal spatial integration for target reconstruction. Training solves a constrained optimization problem:
@@
where $\mathcal{C}(\cdot)$ is the loss function. At convergence, FECNet weights $\xi$ form a feature filter $\phi$. Applying $\phi$ gives:
@@
As a linear operator, $\phi$ has the projection property: